rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#2566 — FS#6528 — Networks 46.105.224 and 225

Attached to Project— Network
Incident
Whole Network
CLOSED
100%
One of the switches N5 managing this network is unstable.
We are looking for the origin of the problem.
Date:  Saturday, 05 May 2012, 02:56AM
Reason for closing:  Done
Comment by OVH - Tuesday, 27 March 2012, 02:01AM

Mar 26 20:52:22 sw.178.33.224.248-n5.fr.eu : 2012 Mar 26 19:58:22 CEST: %ETH_PORT_CHANNEL-5-FOP_CHANGED: port-channel100: first operational port changed from Ethernet1/3 to none
Mar 26 20:52:22 sw.178.33.224.248-n5.fr.eu : 2012 Mar 26 19:58:22 CEST: %FEX-5-FEX_PORT_STATUS_NOTI: Uplink-ID 1 of Fex 100 that is connected with Ethernet1/3 changed its status from Active to Disconnected
Mar 26 20:52:22 sw.178.33.224.248-n5.fr.eu : 2012 Mar 26 19:58:22 CEST: %NOHMS-2-NOHMS_ENV_FEX_OFFLINE: FEX-100 Off-line (Serial Number SSI15430NRS)
Mar 26 20:52:22 sw.178.33.224.248-n5.fr.eu : 2012 Mar 26 19:58:22 CEST: dstats: <{eth_dstats}>dstats_handle_lc_offline: dstats_handle_lc_offline: Processing LC_OFFLINE msg: slot - 32
Mar 26 20:52:22 sw.178.33.224.248-n5.fr.eu : 2012 Mar 26 19:58:22 CEST: dstats: <{eth_dstats}>dstats_clear_slot_delta: dstats_clear_slot_delta: entry


Comment by OVH - Tuesday, 27 March 2012, 02:03AM

We fixed errors on several ports.
We are moving them on other ports of the switch.


Comment by OVH - Tuesday, 27 March 2012, 02:03AM

Ethernet1/19 is up
0 runts 0 giants 26 CRC 0 no buffer
Ethernet1/20 is up
0 runts 0 giants 32 CRC 0 no buffer
Ethernet1/21 is up
0 runts 0 giants 70 CRC 0 no buffer


Comment by OVH - Tuesday, 27 March 2012, 02:04AM

Errors remain,we 'll try to change the cables.


Comment by OVH - Tuesday, 27 March 2012, 02:08AM

Still errors. Errors are on the link vPC and on the links to the fex107. On the two switches,we will try to reset the fex 107.
There will be a traffic interruption for the servers connected to it.


Comment by OVH - Tuesday, 27 March 2012, 02:09AM

Errors remain, we are replacing the fex.


Comment by OVH - Tuesday, 27 March 2012, 02:11AM

The switch seems stable, we are monitoring its performance.


Comment by OVH - Wednesday, 28 March 2012, 22:17PM

Regarding Cisco constructor advices, we are expecting the change of the broken switch during the night of Thursday the 29th of March until Friday the 30th of March.


Comment by OVH - Thursday, 29 March 2012, 23:26PM

We didn't notice problems since 2 days. We reported the maintenance to the next week.